1. Consider the following simplified financial statements for York Corporation. The company has predicted a sales increase of 15%. It has predicted that every item on the balance sheet will increase by 15% as well. Create the pro forma statements = and reconcile them

Income Statement   

Sales $36,000   
Costs of goods $29,800
Net Income $6,200

Balance Sheet

Assets $26,400 Debt $6,300
Equity $20,100
Total Assets $26,400 Total Liab's & Equity $26,400

2. Assume York Corp. (above) pays out half of the net income in the form of a cash dividend. Costs and assets vary with sales, but debt and equity do not. Prepare the pro forma statements and determine the external financing needed.
Answer #1
Income Statement
Projected = A*B
A B
Amount Percentage increase (15%) Amount
Sales 36000 115% 41400
cost of goods sold 29800 115% 34270
Net income 6200 115% 7130
Dividend 3100 115% 3565
Retained Earnings 3100 115% 3565
Balance sheet
A B Projected = A*B Projected = C*D
Amount Percentage increase (15%) Amount Amount Percentage increase (15%) Amount
Debt 6300 115% 7245
Assets 26400 115% 30360 Equity 20100 115% 23115
Total 26400 30360 Total 26400 30360
External Finance need =Current Total Asset or Current Total Liability - Projected Total Asset or Total Liability = 30360-26400 3960